OpenStudy (anonymous):

Questions:- (i)Identify the sequence 3,-3/2,3/4,-3/8,... . (ii)Write the 5th term of the G.P 5,2,4/5,... . (iii)Which term of the A.P 1,5,9,... is 81?

OpenStudy (anonymous):

ANSWERS = (i)G.P (ii)T5=16/125 (iii)n=21

OpenStudy (anonymous):

so the 1st part is G.P

OpenStudy (anonymous):

What should i do in the 2nd one?

OpenStudy (anonymous):

b.Here,Common ratio is 2/5 to find 5th term T5=a*r^(5-1)

OpenStudy (anonymous):

and whats the formula for 3rd part

OpenStudy (anonymous):

t=a+(n-1)d

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

here t=81,a=1,n=?,d=4